naming the slope and point on the line given a point-slope equation
y+3=6(x-2)

Answer by Cromlix(4381)

First put into y = mx + c form by multiplying
and sorting out
y + 3 = 6(x - 2)
y + 3 = 6x - 12
y = 6x - 12 - 3
y = 6x - 15
Secondly pull out of the equation
m - the gradient(slope) = 6
Thirdly make x in the equation = 0
(this is where the line cuts the y axis)
y = 6(0) - 15
so y = - 15 (this is the c in the equation)
Coordinates (0. -15)
Hope this helps
:-)